# Do Test Scores Help Teachers Give Better Track Advice to Students?

### v1

@@ 1 Introduction
Every year, more than one million EU students, upon completing primary education, choose a secondary school track based on recommendations received from their teachers. In some institutional settings—such as the Dutch school system examined in this paper—students may begin secondary education in a more demanding track than the one recommended to them, only under exceptional circumstances. Despite the potentially dramatic consequences that bad track recommendations may have, there is no clear evidence that assigning to teachers alone the role of providing these recommendations is the best solution and generates an optimal track assignment. [^1] There is a wide literature on teachers’ biases that may affect the quality of secondary school track recommendations. Recent examples are: Driessen et al. (2008), Burgess and Greaves (2013), Gerhenson et al. (2016), Osikominu et al. (2021), van Leest et al. (2021), Geven et al. (2021), Carlana and Fort (2022), Carlana et al. (2022a), Carlana et al. (2022b), Ferman and Fontes (2022), Batruch et al. (2023) Bach (2023), Alesina et al. (2024), van Huizen et al. (2024), and Falk et al. (2025). Our goal in this paper is to establish, using a well-defined metric and a quasi-randomized experiment, whether teachers can provide “better” recommendations if they do so after seeing the scores of standardized tests taken by their students.
@@
Let’s divide the population of students into groups – principal strata (Frangakis and Rubin, 2002) – that differ by how the education outcome depends on track recommendations. The first group in which we are interested comprises students for whom the initial track recommendation determines unequivocally the completed track: if the student is recommended the low track in the first year, she will graduate in the same track at the end of secondary school, while if recommended the high track, this will be her final graduation outcome. These are students for whom the teacher’s recommendation is a self-fulfilling prophecy, thus becoming crucially important. In light of this consideration, we call these subjects “Helpable” (H) to signal that they are the ones who can be helped by a high first-year track recommendation instead of a low one. [^5] They correspond to the group of cases that, in the context of Imai et al. (2023), are labeled as “Preventable” crimes. Similarly, de Ree et al. (2025) label them as “Trapped-in-Track” (TT). Adapting the methodology proposed by Imai et al. (2023) to this context, we can establish whether the information provided by test scores helps teachers give better recommendations by measuring the fraction of students in the H group that are recommended the high track by teachers and see if it increases when teachers can decide based on test score information. In the Principal Stratification framework, this is the Average Principal Causal Effect ($APCE$) of “changing the source” of a recommendation (i.e., giving test score information to the source) within the principal stratum of H students. If this $APCE$ is positive, test scores help teachers provide better track advice by recommending more challenging tracks to students who can effectively complete them.
@@
Given the opposite desirability of the findings concerning the effect of test scores on the quality of teacher advice for H and AH students on one hand and AL students on the other, it is natural to seek a method to obtain a comprehensive evaluation. A possible approach is offered by the classification framework proposed by Ben-Michael et al. (2024). This framework enables us to construct a weighted sum of the losses generated by mistakes in recommendations for the various types of students. We can then test whether this sum decreases when teachers see test scores, depending on the relative weight of the short-term losses suffered by students who must change track because of misplacement versus the lifetime losses of students who are not directed to a more challenging track they can eventually complete. We find that the relative weight of short-term losses would have to be unreasonably high to conclude that teachers should not be allowed to upgrade based on test scores.
Finally, building on Imai and Jiang (2022), a corollary contribution of this analysis is that it makes it possible to evaluate whether being allowed to upgrade based on test scores helps teachers improve the fairness of their recommendations with respect to protected attributes such as SES, race, or gender. For example, suppose that gender is a protected attribute with respect to which we want to assess how fair a recommender is. Then, in the stratum of H students, the recommender is fair if the fraction of high recommendations is the same for female and male students. However, the analogous fraction in other strata could be lower or higher as long as it is equal for both genders. Therefore, in the overall population, females may receive a different fraction of high first-year advice from this recommender; however, this would not constitute a violation of fairness, as it would be due to differences in the gender composition of the strata, not to discrimination based on gender within a stratum. We find that the availability of test scores improves the fairness of teachers’ recommendations. Specifically, it raises the chances that immigrants and low-SES students, who are otherwise discriminated against, are recommended a more difficult track that they can complete relative to natives and high-SES peers, respectively.
@@ 2 The Dutch school system and the quasi-experiment
Track
Students initially
assigned to the track
denoted by the row
Test score cut-
off for upgrade
to next level
Fraction above the cutoff
of students eligible
for upgrade to next level
@@
Using the database of the Dutch National Cohort Study on Education (NCO), we consider the population of students who begin their secondary school in the academic years 2015–16 and 2016–17 (hereafter, 2015 and 2016 cohorts). [^7] This database can be accessed following the instructions provided at this link. For documentation in English, see here). The track assignment regulations that generate the quasi-experiment we exploit became effective for students starting secondary school in the 2014-15 academic year, but the data for this cohort do not contain all the necessary information and thus cannot be used. The timing implied by these regulations is described in Figure 1. Let $t=1$ denote the first year of secondary school. In March of the previous year, $t=0$, the primary education teachers give their students an initial track recommendation. The first and second column of Table 1 report the number of students initially assigned to each track by their primary education teachers as a result of this provisional recommendation.
Figure 1: Time-line for track assignment and outcomes
@@
Next, we denote with $Y_{i}\in\{0,1\}$ the outcome that we study. We define $Y_{i}=1$ if student $i$ graduates from a higher track than the one provisionally recommended to him/her in March of $t=0$, before taking the CITO test, and 0 otherwise. We consider a student as graduating in a given track even he/she does so in more years than the required minimum. [^11] Our conclusions do not change qualitatively if we assume instead that a student graduates in a given track only if he/she does so in the regular possible time (results available from the authors). Four years are typically required for a vocational track, while five and six years are needed for the two academic tracks (HAVO and VWO, respectively). Six years are also needed for students who change from VMBO to HAVO, and seven for those who change from HAVO to VWO. The release of the NCO data at our disposal follows the 2015 cohort for 7 years, and the 2016 one for only 6 years after enrollment. Graduation is not observed for some students in these cohorts. We thus assume that the students of these cohorts who are still enrolled in secondary school in the last year of observation will graduate in the track in which they are enrolled. Our goal is to evaluate which source of track recommendation (“teacher alone” or “teacher informed by the test”) is “preferable” according to two criteria: a recommendation is better if it (i) induces a student to enroll in the most difficult track that she/he is able to complete successfully and (ii) reduces to a minimum the chances of costly track changes.
@@ 3.2 Students’ strata
Notes: The four panels refer to the four Principal Strata listed in the first column of the table and defined by the possible values of $R_{i}$ and $Y_{i}$: the Always-Low (AL), the Always-High (AH), the Helpable (H) and the Rebels (R). Columns 2 and 3 describe the values of the potential outcomes $Y_{i}$ when the treatment $R_{i}$ is equal to 0 or $1,$, respectively, and the Exclusion Restriction assumption holds. Each one of these Principal Strata is divided into additional strata defined by the possible values of $Z_{i}$ and $R_{i}$, as listed in the fourth column: the Never Takers (NT), the Always Takers (AT), the Compliers (C) and the Defiers (D). Columns 5 and 6 describe the values of the potential treatments $R_{i}$ when the assignment to treatment $Z_{i}$ is equal to 0 or $1,$, respectively. The Notes in the first three rows of the last column make clear that Defiers may or may not exist depending on whether the Monotonicity Assumption 5 of $R_{i}$ with respect to $Z_{i}$ holds. The last note in the fourth row states that given the maintained Monotonicity Assumption 4 of $Y_{i}$ with respect to $R_{i}$, the Rebels do not exist.
@@ 3.3 The metric to compare recommendations
The first criterion defines what constitutes a better recommendation for Helpable (H) students. In their case, the goal of teachers should be to recommend the high track to the largest fraction of these students because, as a result of this recommendation, they will complete precisely this track, achieving a better outcome for themselves as well as for society. Therefore, we need to estimate the difference between “the fraction of H students who are recommended a high first-year track by teachers who have seen students’ test scores” and “the analogous fraction for teachers who have not seen them”. If this difference is positive, test scores help teachers give students better track advice.
@@
The $APCE_{J}$ is the difference between the fraction of students in stratum $J$ who are recommended the high track by the teachers informed by test scores ($Z_{i}=1$) minus the fraction receiving the same recommendation by teachers who have not seen the test scores ($Z_{i}=0$). To put it differently, the $APCE_{J}$ is the Intention To Treat effect of the source of recommendations on their content in stratum $J$.
@@
If Monotonicity of $R_{i}$ with respect to $Z_{i}$ (Assumption 5) holds, Defiers do not exist, and the three $APCE_{J}$ are all non-negative by construction, being equal to the proportion of Compliers. In this case we could conclude [^17] As already mentioned when we introduced Assumption 5 in the previous section, we emphasize that this assumption allows to interpret $APCE_{J}$ as the proportion of Compliers in stratum $j$, but it is not necessary for identification. that test scores help teachers give better recommendations if the $APCE_{H}$ and the $APCE_{AH}$ are positive while the $APCE_{AL}$ is equal to zero. Having established that the $APCE_{J}$ are informative about whether test scores help teachers give better recommendations, in the next section, we show how these population parameters can be identified with data generated by the quasi-experiment that we study.
@@ 3.4 Identification of the $APCE_{J}$
where the terms in the numerator do not depend on missing potential outcomes and can be estimated with their sample analogs. The denominator is instead the proportion of Helpable students, given that Rebels are excluded because of the Monotonicity Assumption 4, and can only be partially identified. In some settings, even if principal strata are latent, their proportions can be point identified under specific assumptions. This is the case, for example, of the proportion of compliers in IV settings. In our case, because the principal strata are defined by the joint values of $Y(r)$, the strata proportions are not point identified unless we impose further assumptions on the distribution of potential outcomes and the recommendation $R$.
@@ 4.1 Estimates for the three strata
Focusing for brevity just on Columns 4 and 5 of the table, the gap between the lower and the upper bounds of the $APCE_{AH}$ is remarkably small in almost all tracks, allowing for a very precise estimation of this population parameter. [^22] In the case of HAVO track, the sample analog of the numerator in the RHS of equation (8) is close to zero but has a negative sign, possibly due to small sample variability. We approximate it with 0. The intuition for why we obtain such a precise result is that we can be confident in the population proportion of AH students (i.e. the denominator) as those with $R=0$ and $Y=1$ are for sure AH, and measuring this under the cutoff (i.e. for those with $Z=0$) gives a precise estimate of the population parameter, showing that the population proportion of AH is approximately 26%. [^23] More specifically, because it is extremely rare to have $Z=0$ and $R=1$ (which can only occur due to the institutional complications explained in footnote 12) we know that among those with $Z=0$, those with $R=0$ and $Y=1$ are AH for sure, while those with $R=0$ and $Y=0$ are for sure not AH, and these two categories together form virtually the whole population. This means that the fraction of AH students can be estimated very precisely, as $Z$ is randomized and the same fraction should exist above the threshold. In the eight basic tracks, the point estimates of the $APCE$ range from 0 to 13%. When the tracks are aggregated in the Vocational, Academic, and All tracks (last three rows of the table), the $APCE_{AH}$ is estimated to be about 6.2%, 4.3%, and 5.6%, respectively, and statistically significant. These results indicate that when teachers receive test score information, they send a significantly higher fraction of AH students to the high track, thereby reducing the number of costly track changes this group would otherwise incur.
@@
Focusing again for brevity just on the last two columns of the table, Column 4 shows that the estimated lower bound of the $APCE_{AL}$ is positive and statistically different from zero in many tracks, ranging from 4% for BL to 14% for HAVO/VWO. Considering the aggregate tracks in the last three rows of the table, the $APCE_{AL}$ is at least as high as 7% and significantly different from zero in all of them. Therefore, it appears that the information provided by test scores leads teachers to recommend the high track to too many AL students. The consequence is a considerable deterioration of the quality of their recommendations. A possible reason for this finding is that primary school teachers do not bear any cost deriving from secondary school track changes and are therefore less sensitive to this consequence of the advice they offer to their students.
@@ 4.2 A comprehensive evaluation
The evidence in the previous section suggests that, in the case of Helpable and Always High students, test score information helps teachers improve the quality of the recommendations in terms of the two criteria that we have adopted (directing students towards the most difficult track they can successfully complete and reducing track changes). On the contrary, in the case of the Always Low, the positive $APCE_{AL}$ estimate indicates that the same information reduces this quality because it induces too many AL students to enroll in the high track that they later cannot complete.
@@
where $\Pi_{01}(z)=Pr(R(z)=0;Y(1)=1)$ is the probability of a False Negative (H or AH) and $\Pi_{10}(z)=Pr(R(z)=1;Y(1)=0)$ is the probability of a False Positive (AL). Our goal is to compare the overall loss $\mathcal{L}(l_{10},z)$ when the source is a “teacher alone” $(z=0)$ or a “teacher informed by test scores”$(z=1)$, computed at different values of the relative weight $l_{10}$. In light of the above considerations, this weight is at most equal to $1$, when the short-run loss of AL students is considered by the policy-maker as relevant as the long-run loss of H students, so that $l_{10}=1$, and decreases towards 0 as the former loss becomes less relevant $(l_{10}<1)$.
@@
Failure to reject the null hypothesis of this test (high P-value) indicates that we cannot rule out the possibility that informing teachers with test scores is a better (or equally effective) system because we do not have enough evidence in the data to conclude that the provision of test scores leads to larger losses. Conversely, if we can reject $H_{0}$ (small P-value), we can rule out that teachers informed by test scores generate smaller losses and thus give better recommendations than teachers alone.
With the exception of three tracks (V:BL, V:GT and A:HAVO), we cannot reject that the difference is negative even when the short-run loss of AL students is considered as relevant as the long-run loss suffered by H students $(l_{10}=1)$ and AL students have the maximum relative weight. For the remaining V:BL, V:GT and A:HAVO tracks, we reject that the difference is negative only when the short-term loss suffered by AL students has an (implausibly) high relative weight – higher than about 0.8, 0.5, and 0.3, respectively, at the 5% significance level. Similar evidence is reported in the right panel of the figure for the three aggregate tracks. We never reject the null with the same confidence for the All and Vocational tracks, even if the relative weight on the short-term loss of AL students is the highest. In the case of the Academic track, we reject only when this weight is higher than about 0.75.

### v2

@@ 1 Introduction
Every year, more than one million EU students, upon completing primary education, choose a secondary school track based on recommendations received from their teachers. In some institutional settings—such as the Dutch school system examined in this paper—only under exceptional circumstances students may begin secondary education in a more demanding track than the one recommended to them. Despite the potentially dramatic consequences that bad track recommendations may have, there is no clear evidence that assigning to teachers alone the role of providing these recommendations is the best solution and generates an optimal track assignment. [^1] There is a wide literature on teachers’ biases that may affect the quality of secondary school track recommendations. Recent examples are: Driessen et al. (2008), Burgess and Greaves (2013), Gerhenson et al. (2016), Osikominu et al. (2021), van Leest et al. (2021), Geven et al. (2021), Carlana and Fort (2022), Carlana et al. (2022a), Carlana et al. (2022b), Ferman and Fontes (2022), Batruch et al. (2023) Bach (2023), Alesina et al. (2024), van Huizen et al. (2024), and Falk et al. (2025). Our goal in this paper is to establish, using a well-defined metric and a quasi-randomized experiment, whether teachers can “improve” the quality of their recommendations after seeing the scores of standardized tests taken by their students.
@@
Let’s divide the population of students into groups – principal strata (Frangakis and Rubin, 2002) – that differ by how the education outcome depends on track recommendations. The first group in which we are interested comprises students for whom the track recommendation determines unequivocally the completed track: if the student is recommended the low track in the first year, she will graduate in the same track at the end of secondary school, while if recommended the high track, this will be her final graduation outcome. These are students for whom the teacher’s recommendation is a self-fulfilling prophecy, thus becoming crucially important. In light of this consideration, we call these subjects “Helpable” (H) to signal that they are the ones who can be helped by a high first-year track recommendation instead of a low one. [^5] They correspond to the group of cases that, in the context of Imai et al. (2023), are labeled as “Preventable” crimes. Similarly, de Ree et al. (2025) label them as “Trapped-in-Track” (TT). Adapting the methodology proposed by Imai et al. (2023) to this context, we can establish whether the information provided by test scores helps teachers improve their recommendations by measuring the fraction of students in the H group that are recommended the high track by teachers and see if it increases when teachers can decide to upgrade based on test score information. In the Principal Stratification framework, this is the Average Principal Causal Effect ($APCE$) of “changing the source” of a recommendation (i.e., giving test score information to the source, in our context) within the principal stratum of H students. If this $APCE$ is positive, test scores help teachers provide better track advice by upgrading to more challenging tracks students who can effectively complete them.
@@
Given the opposite desirability of the findings concerning the effect of test scores on the quality of teacher advice for H and AH students on one hand and AL students on the other, it is natural to seek a method to obtain a comprehensive evaluation. An approach is offered by the classification framework proposed by Ben-Michael et al. (2024). This framework enables us to construct a weighted sum of the losses generated by mistakes in recommendations for the various types of students. We can then test whether this sum decreases when teachers can upgrade based on test scores, depending on the relative weight of the short-term losses suffered by students who must change track because of misplacement versus the lifetime losses of students who are not directed to a more challenging track they can eventually complete. We find that the relative weight of short-term losses would have to be unreasonably high to conclude that teachers should not be allowed to upgrade based on test scores.
Finally, building on Imai and Jiang (2022), a corollary contribution of this analysis is that it makes it possible to evaluate whether being allowed to upgrade based on test scores helps teachers improve the fairness of their recommendations with respect to protected attributes such as SES, race, or gender. For example, suppose that gender is a protected attribute with respect to which we want to assess how fair a recommender is. Then, in the stratum of H students, the recommender is fair if the fraction of high recommendations is the same for female and male students. However, the analogous fraction in other strata could be lower or higher as long as it is equal for both genders. Therefore, in the overall population, females may receive a different fraction of high first-year advice from this recommender; however, this would not constitute a violation of fairness, as it would be due to differences in the gender composition of the strata, not to discrimination based on gender within a stratum. We find that allowing teachers to upgrade recommendations based on test scores improves fairness. Specifically, it raises the chances that immigrants and low-SES students, who are otherwise discriminated against, are recommended a more difficult track that they can complete relative to natives and high-SES peers, respectively.
@@ 2 The Dutch school system and the quasi-experiment
Track Students initially assigned to the track denoted by the row Test score cut- off for upgrade to next level Fraction above the cutoff of students eligible for upgrade to next level
@@
Using the database of the Dutch National Cohort Study on Education (NCO), we consider the population of students who begin their secondary school in the academic years 2015–16 and 2016–17 (hereafter, 2015 and 2016 cohorts). [^7] This database can be accessed following the instructions provided at this link. For documentation in English, see here). The track assignment regulations that generate the quasi-experiment we exploit became effective for students starting secondary school in the 2014-15 academic year, but the data for this cohort do not contain all the necessary information and thus cannot be used. The timing implied by these regulations is described in Figure 1. Let $t=1$ denote the first year of secondary school. In March of the previous year, $t=0$, the primary education teachers give their students an initial track recommendation. The first and second columns of Table 1 report the number of students in each cohort initially assigned to each track by their primary education teachers as a result of this provisional recommendation.
Figure 1: Time-line for track recommendations and outcomes
@@
Next, we denote with $Y_{i}\in\{0,1\}$ the outcome that we study. We define $Y_{i}=1$ if student $i$ graduates from a higher track than the one provisionally recommended to him/her in March of $t=0$, before taking the CITO test, and 0 otherwise. We consider a student as graduating in a given track even if he/she does so in more years than the required minimum. [^11] Our conclusions do not change qualitatively if we assume instead that a student graduates in a given track only if he/she does so in the regular possible time (results available from the authors). Four years are typically required for a vocational track, while five and six years are needed for the two academic tracks (HAVO and VWO, respectively). Six years are also needed for students who change from VMBO to HAVO, and seven for those who change from HAVO to VWO. The release of the NCO data at our disposal follows the 2015 cohort for 7 years, and the 2016 one for only 6 years after enrollment. Graduation is not observed for some students in these cohorts. We thus assume that the students of these cohorts who are still enrolled in secondary school in the last year of observation will graduate in the track in which they are enrolled. Our goal is to evaluate which source of track recommendation (“teacher alone” or “teacher informed by test scores and allowed to upgrade”) is “preferable” according to two criteria: a recommendation is better if it (i) induces a student to enroll in the most difficult track that she/he is able to complete successfully and (ii) reduces to a minimum the chances of costly track changes.
@@ 3.2 Students’ strata
Notes: The four panels refer to the four Principal Strata listed in the first column of the table and defined by the possible values of $R_{i}$ and $Y_{i}$: the Always-Low (AL), the Always-High (AH), the Helpable (H) and the Rebels (R). Columns 2 and 3 describe the values of the potential outcomes $Y_{i}$ when the treatment $R_{i}$ is equal to 0 or $1$, respectively, and the Exclusion Restriction assumption holds. Each one of these Principal Strata is divided into additional strata defined by the possible values of $Z_{i}$ and $R_{i}$, as listed in the fourth column: the Never Takers (NT), the Always Takers (AT), the Compliers (C) and the Defiers (D). Columns 5 and 6 describe the values of the potential treatments $R_{i}$ when the assignment to treatment $Z_{i}$ is equal to 0 or $1$, respectively. The Notes in the first three rows of the last column make clear that Defiers may or may not exist depending on whether the Monotonicity Assumption 5 of $R_{i}$ with respect to $Z_{i}$ holds. The last note in the fourth row states that given the maintained Monotonicity Assumption 4 of $Y_{i}$ with respect to $R_{i}$, the Rebels do not exist.
@@ 3.3 The metric to compare recommendations
The first criterion defines what constitutes a better recommendation for Helpable (H) students. In their case, the goal of teachers should be to recommend the high track to the largest fraction of these students because, as a result of this recommendation, they will complete precisely this track, achieving a better outcome for themselves as well as for society. Therefore, we need to estimate the difference between “the fraction of H students who are recommended a high first-year track by teachers who can upgrade based on students’ test scores” and “the analogous fraction when teachers do not see test scores”. If this difference is positive, test scores help teachers improve the track advising they provide to their students.
@@
The $APCE_{J}$ is the difference between the fraction of students in stratum $J$ who are recommended the high track by teachers who can upgrade based on test scores ($Z_{i}=1$) minus the fraction receiving the same recommendation by teachers who have not seen the test scores ($Z_{i}=0$). To put it differently, the $APCE_{J}$ is the Intention To Treat effect of the source of recommendations on their content in stratum $J$.
@@
If Monotonicity of $R_{i}$ with respect to $Z_{i}$ (Assumption 5) holds, Defiers do not exist, and the three $APCE_{J}$ are all non-negative by construction, being equal to the proportion of Compliers. In this case we could conclude [^17] As already mentioned when we introduced Assumption 5 in the previous section, we emphasize that this assumption allows to interpret $APCE_{J}$ as the proportion of Compliers in stratum $j$, but it is not necessary for identification. that test scores help teachers improve their recommendations if the $APCE_{H}$ and the $APCE_{AH}$ are positive while the $APCE_{AL}$ is equal to zero. Having established that the $APCE_{J}$ are informative about whether test scores help teachers improve the quality of track advising, the next section shows how these population parameters can be identified using data generated by the quasi-experiment under study.
@@ 3.4 Identification of the $APCE_{J}$
where the terms in the numerator do not depend on missing potential outcomes and can be estimated with their sample analogs. The denominator is instead the proportion of Helpable students, given that Rebels are excluded because of the Monotonicity Assumption 4, and can only be partially identified. In some settings, even if principal strata are latent, their proportions can be point identified under specific assumptions. This is the case, for example, of the proportion of compliers in IV settings, because $Z$ is random. In our case, because the principal strata are defined by the joint values of $Y(r)$ and $R$ is confounded, the strata proportions are not point-identified unless we impose further assumptions on the distribution of potential outcomes and the recommendation $R$.
@@ 4.1 Estimates for the three strata
Focusing for brevity just on Columns 4 and 5 of the table, the gap between the lower and the upper bounds of the $APCE_{AH}$ is remarkably small in almost all tracks, allowing for a very precise estimation of this population parameter. [^22] In the case of HAVO track, the sample analog of the numerator in the RHS of equation (8) is close to zero but has a negative sign, possibly due to small sample variability. We approximate it with 0. The intuition for why we obtain such a precise result is that we can be confident in the population proportion of AH students (i.e. the denominator) as those with $R=0$ and $Y=1$ are for sure AH, and measuring this under the cutoff (i.e. for those with $Z=0$) gives a precise estimate of the population parameter, showing that the population proportion of AH is approximately 26%. [^23] More specifically, because it is extremely rare to have $Z=0$ and $R=1$ (which can only occur due to the institutional complications explained in footnote 12) we know that among those with $Z=0$, those with $R=0$ and $Y=1$ are AH for sure, while those with $R=0$ and $Y=0$ are for sure not AH, and these two categories together form virtually the whole population. This means that the fraction of AH students can be estimated very precisely, as $Z$ is randomized and the same fraction should exist above the threshold. In the eight basic tracks, the point estimates of the $APCE$ range from 0 to 13%. When the tracks are aggregated in the Vocational, Academic, and All tracks (last three rows of the table), the $APCE_{AH}$ is estimated to be about 6.2%, 4.3%, and 5.6%, respectively, and statistically significant. These results indicate that when teachers are allowed to upgrade recommendations based on test scores, they send a significantly higher fraction of AH students to the high track, thereby reducing the costly track changes this group would otherwise incur.
@@
Focusing again for brevity just on the last two columns of the table, Column 4 shows that the estimated lower bound of the $APCE_{AL}$ is positive and statistically different from zero in many tracks, ranging from 4% for BL to 14% for HAVO/VWO. Considering the aggregate tracks in the last three rows of the table, the $APCE_{AL}$ is at least as high as 7% and significantly different from zero in all of them. Therefore, it appears that the information provided by test scores leads teachers to upgrade AL students into the high track. The consequence is a deterioration of the quality of recommendations because AL students cannot complete the high track. A possible reason for this finding is that primary school teachers do not incur any costs resulting from secondary school track changes and are therefore less sensitive to these consequences of the advice they offer to their students.
@@ 4.2 A comprehensive evaluation
The evidence in the previous section suggests that, in the case of Helpable and Always High students, test score information helps teachers improve the quality of their recommendations in terms of the two criteria that we have adopted (directing students towards the most difficult track they can successfully complete and reducing track changes). On the contrary, in the case of the Always Low, the positive $APCE_{AL}$ estimate indicates that the same information reduces this quality because it induces AL students to enroll in the high track that they later cannot complete.
@@
where $\Pi_{01}(z)=Pr(R(z)=0;Y(1)=1)$ is the probability of a False Negative (H or AH) and $\Pi_{10}(z)=Pr(R(z)=1;Y(1)=0)$ is the probability of a False Positive (AL). Our goal is to compare the overall loss $\mathcal{L}(l_{10},z)$ when the source is a “teacher alone” $(z=0)$ or a “teacher informed by test scores and allowed to upgrade”” $(z=1)$, computed at different values of the relative weight $l_{10}$. This weight is at most equal to $1$, when the short-run loss of AL students is considered by the policy-maker as relevant as the long-run loss of H students, so that $l_{10}=1$, and decreases towards 0 as the former loss becomes less relevant $(l_{10}<1)$.
@@
Failure to reject the null hypothesis of this test (high P-value) indicates that we cannot rule out the possibility that allowing teachers to upgrade their recommendations based on test scores is a better (or equally effective) system because we do not have enough evidence to conclude that the provision of test scores leads to larger losses. Conversely, if we can reject $H_{0}$ (with a small P-value), we can rule out the possibility that allowing teachers to upgrade based on test scores generates smaller losses and thus improves recommendations.
With the exception of three tracks (V:BL, V:GT and A:HAVO), we cannot reject that the difference is negative even when the short-run loss of AL students is considered as relevant as the long-run loss suffered by H students $(l_{10}=1)$ and AL students have the maximum relative weight. For the remaining V:BL, V:GT and A:HAVO tracks, we reject that the difference is negative only when the short-term loss suffered by AL students has an (implausibly) high relative weight – higher than about 0.8, 0.5, and 0.3, respectively, at the 5% significance level. Similar evidence is reported in the right panel of the figure for the three aggregate tracks. We never reject the null at the same significance level for the All and Vocational tracks, even if the relative weight on the short-term loss of AL students is the highest. In the case of the Academic track, we reject only when this weight is higher than about 0.75.